Compound X contains only carbon, hydrogen and oxygen. A 0.2000 g sample of compound X was burnt completely in a combustion vessel to yield 0.3906 g of C02 and 0.200 g of H20. Determine the empirical formula of compound X.


1
Expert's answer
2022-04-14T08:26:03-0400

"C_xH_yO_z +O_2 \rightarrow CO_2 + H_2O"

M(CO2)=12+16x2=44 g/mol

M(H2O)=2x1+16=18 g/mol

n(CO2)=0.3906/44=0.00888 mol

n(C)=n(CO2)=0.00888 mol

n(OCO2)=2n(CO2)=0.0178 mol

n(H2O)=0.2/18=0.0111 mol

n(H)=2n(H2O)=0.0222 mol

n(OH2O)=n(H2O)=0.0111:mol

n(O)=n(O)CO2+n(O)H2O=0.0178+0.0111=0.0289 mol

Let delete all on n(C)

C-1

H-2.5

O-3.25

Then substance is "C_xH_{2.5x}O_{3.25x}"

Formula is "C_4H_{10}O_{13}"